Verify that each of the following statements is a false statement by finding a counterexample for each.

  1. xx=1
  2. x+33=x+1
  3. x2+16=x+4
1
Expert's answer
2022-05-11T08:59:23-0400

1. Let x=0, then "0(0) \not =1"

So xx=1 is false

2. Let x=1, then "1+33 \not= 1+1 \\
34 \not=2"

So x+33=x+1 is false


3. Let x=1

"1^2+16 \not=1+4\\
17 \not=5"

So x2+16=x+4 is false.